OpenStudy (anonymous):

How do I Calculate the distance between the points =G−8,8 and =F−5,2 in a Cartesian plane. I must write the exact answer in radical form.

OpenStudy (anonymous):

You may use the distance formula \[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}} \]

OpenStudy (anonymous):

(5-8)^2+(2-8)^2?

OpenStudy (anonymous):

Your points are (G-8,8) and (F-5,2) so your x1 would be G-8

OpenStudy (anonymous):

same goes for x2

OpenStudy (anonymous):

thats what I had do I change it to positive ?

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Unless you meant G = (-8,8)

OpenStudy (anonymous):

it is (-8,8)

OpenStudy (anonymous):

I'm going off with what you gave me as your question is not very clear.

OpenStudy (anonymous):

Thought so. Ok so \[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}} \implies \sqrt{(-5-8)^2+(2-8)^2}\]

OpenStudy (anonymous):

As your points are G = (-8,8) an F = (-5,2) :) does that make sense?

OpenStudy (anonymous):

Yes i thought i typed that

OpenStudy (anonymous):

You may do the simplification now, remember it's squared so your negative becomes positive.

OpenStudy (anonymous):

Oh notice there's an error, it should be x = +5 and y = -2 so what you wrote is right, I was still using what you gave me.

OpenStudy (anonymous):

radical 205

OpenStudy (anonymous):

So just to put the final using what your picture tells us, this is what you should have: \[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}} \implies \sqrt{(5-8)^2+(-2+8)^2}\]

OpenStudy (anonymous):

\[\sqrt{45}\]

OpenStudy (anonymous):

That's good :), and sorry about earlier, it's very frustrating when you keep losing connection and then the information is everywhere.

OpenStudy (anonymous):

Is that in simplest form?

OpenStudy (anonymous):

Yes, leave it as radical :)

OpenStudy (anonymous):

Or I guess you could put it as \[3 \sqrt{5}\]

OpenStudy (anonymous):

Yes I believe \[3 \sqrt{5}\] is what they might want :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Np! Take care.